# Analysis of neutrino oscillation parameters in the light on quantum entanglement

### v1

@@ 4.3 Oscillation parameters at entanglement extrema
In the previous subsections, we identified three distinct energy windows where EOF and concurrence exhibit “global minimum”, maximum, and “local minimum” values. To summarize, EW1 shows $\nu_{\mu}$ in an almost pure state, being completely disentangled from $\nu_{e}$ and $\nu_{\tau}$, EW2 exhibits maximal entanglement among all three flavor states and in EW3, there is minimal mixing between $\nu_{\mu}$ and $\nu_{\tau}$, with $\nu_{e}$ remaining nearly separable. From an experimental perspective, precise measurements of neutrino oscillation parameters are most effectively conducted in energy windows where one of the three flavor states is either in a pure state or nearly disentangled. Specifically, EW1 provides a pure state for $\nu_{\mu}$, while EW3 offers a nearly separable state for $\nu_{e}$. In this subsection, we will focus on the energy points of 1.27 GeV and 2.5 GeV to determine the best-fit values for the oscillation parameters at these specific points. To achieve this, we varied all five oscillation parameters ($\theta_{12},\theta_{13},\theta_{23}$ $,\Delta m_{21}^{2}$, and $\Delta m_{31}^{2}$) within their $3\sigma$ confidence levels as specified in NuFit v5.2 Esteban:2020cvm , while considering the full range of $\delta_{CP}$ from $0^{\circ}$ to $360^{\circ}$. We performed a comprehensive analysis by scanning the oscillation parameters and identified the minimum concurrence value along with the associated parameters. This analysis was performed using the DUNE experimental configuration.
@@
The results are summarized in Table 2. The first row presents the best-fit values of oscillation parameters from NuFit v5.2. The second and third rows show the oscillation parameters at 2.5 GeV and 1.27 GeV, respectively, where the concurrence is at its minimum, with the minimum concurrence values noted in parentheses. Table 2 reveals notable differences between the oscillation parameters at these energy points and the values from NuFit v5.2. For instance, while $\theta_{12}$ is $33.41^{\circ}$ in NuFit v5.2, it shifts to $31.31^{\circ}$ and $35.71^{\circ}$ at 2.5 GeV and 1.27 GeV, respectively. Interestingly, $\theta_{13}$ values remain relatively consistent across these reference cases. However, the octant problem of the atmospheric angle shows that $\theta_{23}$, which is in the lower octant according to NuFit v5.2, shifts towards a maximally mixing state at both minimum concurrence points. Additionally, solar and atmospheric mass squared differences also change from the NuFit values. Most notably, $\delta_{CP}$ trends towards the maximum CP-violating phase of $270^{\circ}$ when considering the minimum concurrence points. In the next section, we will explore how the concepts of CP violation, mass hierarchy, and the octant of the atmospheric mixing angle are influenced by the presence or absence of quantum entanglement.
@@ 5 Results
Condition $\theta_{12}$ $\theta_{13}$ $\theta_{23}$ $\Delta m_{21}^{2}$ (${\rm eV}^{2}$) $\Delta m_{31}^{2}$ (${\rm eV}^{2}$) $\delta_{CP}$
Nufit v5.2 Esteban:2020cvm $33.41^{\circ}$ $8.58^{\circ}$ $42.2^{\circ}$ $7.41\times 10^{-5}$ $2.507\times 10^{-3}$ $232^{\circ}$
@@@@ 2.50 GeV (0.03225) $31.31^{\circ}$ $8.43^{\circ}$ $45.2^{\circ}$ $7.01\times 10^{-5}$ $2.487\times 10^{-3}$ $270^{\circ}$
@@@@ 1.27 GeV (0.00314) $35.71^{\circ}$ $8.23^{\circ}$ $45.2^{\circ}$ $7.61\times 10^{-5}$ $2.587\times 10^{-3}$ $270^{\circ}$
Table 2: First row presents the values of the oscillation parameters from NuFit v5.2, while the second (third) row shows the best-fit values obtained by considering the “local (global) minima” of concurrence. The corresponding concurrence values are indicated in parentheses in the first column of the second and third rows.
@@ 5.1 CP violation
The CP violation sensitivity results are illustrated in the left panel of Fig. 4. In this panel, the red, blue, and green curves represent the CP violation sensitivity as a function of the true value of $\delta_{CP}$ under three different conditions: the red curve corresponds to the true values of oscillation parameters from the first row of Table 2, the blue and green curves represent the sensitivity using the best-fit values of oscillation parameters from the second and third rows of Table 2, respectively. In terms of quantum entanglement, if entanglement is not considered, the red curve shows the CP violation sensitivity for the DUNE configuration. However, when entanglement is taken into account, the CP violation sensitivity decreases for both the “global” and “local” minima concurrence points, dropping from $8.5\sigma$ to $8.2\sigma$ and to $7.0\sigma$, respectively.
@@ 6 Concluding remarks
Focusing on the behavior of EOF and concurrence, we identify three energy windows where “global minima”, maxima, and “local minima” occur. For the “local” and “global” minima points, we derive the best-fit oscillation parameters by minimizing entanglement. We extend these results at the minima points to examine sensitivities to CP violation (CPV), mass hierarchy, and the octant of the atmospheric mixing angle. In the presence of quantum entanglement, CPV sensitivity was reduced from $8.5~{}\sigma$ to $8.2~{}\sigma$ and $7.0~{}\sigma$ for the “global” and “local” minima, respectively. In contrast to CPV, the sensitivity to the mass hierarchy increase when considering the concept of entanglement. However, for octant sensitivity, due to the atmospheric mixing angle approaching maximal mixing at the points of minimum entanglement, there is no significant improvement in sensitivity in the presence of entanglement.

### v2

@@ 4.3 Oscillation parameters at entanglement extrema
In the previous subsections, we identified three distinct energy windows where EOF and concurrence exhibit “global minimum”, maximum, and “local minimum” values. To summarize, EW1 shows $\nu_{\mu}$ in an almost pure state, being completely disentangled from $\nu_{e}$ and $\nu_{\tau}$, EW2 exhibits maximal entanglement among all three flavor states and in EW3, there is minimal mixing between $\nu_{\mu}$ and $\nu_{\tau}$, with $\nu_{e}$ remaining nearly separable. From an experimental perspective, precise measurements of neutrino oscillation parameters are most effectively conducted in energy windows where one of the three flavor states is either in a pure state or nearly disentangled. Specifically, EW1 provides a pure state for $\nu_{\mu}$, while EW3 offers a nearly separable state for $\nu_{e}$. In this subsection, we will focus on the energy points of 1.27 GeV and 2.5 GeV to determine the best-fit values for the oscillation parameters at these specific points. The main motivation to focus on these two points are following; at the concurrence minimum, the neutrino flavor state becomes nearly separable, minimizing the parameter uncertainty. Hence, oscillation parameters can be measured with the highest precision. In our analysis, both EOF and concurrence attain absolute minima at 1.27 GeV, where $P_{\mu\mu}$ approaches unity and the appearance probabilities of $\nu_{e}$ and $\nu_{\tau}$ vanish. This signifies a predominantly pure $\nu_{\mu}$ configuration, implying that oscillation parameters measured near this energy are expected to be more precise and reliable. Moreover, the “local minimum” point holds equal significance in our analysis for two main reasons. First, around 2.5 GeV, the entanglement reaches a minimum (though not exactly zero), indicating a nearly separable (but not fully separable) neutrino flavor state, which is particularly suitable for precise extraction of oscillation parameters. Second, the DUNE neutrino flux peaks at approximately 2.5 GeV, providing the highest event rate within this energy range and thereby improving the statistical precision of the results. To achieve this, we varied all five oscillation parameters ($\theta_{12},\theta_{13},\theta_{23}$ $\Delta m_{21}^{2},$ and $\Delta m_{31}^{2}$) within their $3\sigma$ confidence levels as specified in NuFit v6.0 Esteban:2024eli, while considering the full range of $\delta_{CP}$ from $0^{\circ}$ to $360^{\circ}$. We performed a comprehensive analysis by scanning the oscillation parameters and identified the minimum concurrence value along with the associated parameters. This analysis was performed using the DUNE experimental configuration.
@@
The results are summarized in Table 2. The first row presents the best-fit values of oscillation parameters from NuFit v6.0. The second and third rows show the oscillation parameters at 2.5 GeV and 1.27 GeV, respectively, where the concurrence is at its minimum, with the minimum concurrence values noted in parentheses. Table 2 reveals notable differences between the oscillation parameters at these energy points and the values from NuFit v6.0. For instance, while $\theta_{12}$ is $33.68^{\circ}$ in NuFit v6.0, it shifts to $31.31^{\circ}$ and $35.71^{\circ}$ at 2.5 GeV and 1.27 GeV, respectively. Interestingly, $\theta_{13}$ values remain relatively consistent across these reference cases. However, the octant problem of the atmospheric angle shows that $\theta_{23}$, which is in the lower octant according to NuFit v6.0, shifts towards a maximally mixing state at both minimum concurrence points. Additionally, solar and atmospheric mass squared differences also change from the NuFit values. Most notably, $\delta_{CP}$ trends towards the maximum CP-violating phase of $270^{\circ}$ when considering the minimum concurrence points. In the next section, we will explore how the concepts of CP violation, mass hierarchy, and the octant of the atmospheric mixing angle are influenced by the presence or absence of quantum entanglement.
@@ 5 Results
Condition $\theta_{12}(^{\circ})$ $\theta_{13}(^{\circ})$ $\theta_{23}(^{\circ})$ $\Delta m_{21}^{2}$(${\rm eV}^{2}$) $\Delta m_{31}^{2}$(${\rm eV}^{2}$) $\delta_{CP}(^{\circ})$
NuFit v6.0 Esteban:2024eli $33.68^{\circ}$ $8.56^{\circ}$ $43.3^{\circ}$ $7.49\times 10^{-5}$ $2.513\times 10^{-3}$ $212^{\circ}$
@@ 2.50 GeV (0.03225) $31.31^{\circ}$ $8.43^{\circ}$ $45.2^{\circ}$ $7.01\times 10^{-5}$ $2.487\times 10^{-3}$ $270^{\circ}$
@@ 1.27 GeV (0.00314) $35.71^{\circ}$ $8.23^{\circ}$ $45.2^{\circ}$ $7.61\times 10^{-5}$ $2.587\times 10^{-3}$ $270^{\circ}$
Table 2: First row presents the values of the oscillation parameters from NuFit v6.0, while the second (third) row shows the best-fit values obtained by considering the “local (global) minima” of concurrence. The corresponding concurrence values are indicated in parentheses in the first column of the second and third rows.
@@ 5.1 CP violation
The CP violation sensitivity results are illustrated in the left panel of Fig. 5. In this panel, the red, blue, and green curves represent the CP violation sensitivity as a function of the true value of $\delta_{CP}$ under three different conditions: the red curve corresponds to the true values of oscillation parameters from the first row of Table 2, the blue and green curves represent the sensitivity using the best-fit values of oscillation parameters from the second and third rows of Table 2, respectively. In terms of quantum entanglement, if entanglement is not considered, the red curve shows the CP violation sensitivity for the DUNE configuration. When entanglement effects are included, the CP violation sensitivity decreases for both the “global” and “local” minimum concurrence points, reducing from $8.50\sigma$ to $8.25\sigma$ and $7.19\sigma$, respectively, for $\delta_{CP}^{\text{true}}$ values between $0^{\circ}$ and $180^{\circ}$. In the range $\delta_{CP}^{\text{true}}=[180^{\circ}–360^{\circ}]$, the highest CPV sensitivity corresponds to the “global” minimum concurrence point with a value of $7.33\sigma$, while the “local” and NuFIT data points yield maximum sensitivities of $6.37\sigma$ and $7.04\sigma$, respectively.
@@ 6 Concluding remarks
Focusing on the behavior of EOF and concurrence, we identify three energy windows where “global minima”, maxima, and “local minima” occur. For the “local” and “global” minima points, we derive the best-fit oscillation parameters by minimizing entanglement. We extend these results at the minima points to examine sensitivities to CP violation (CPV), mass hierarchy, and the octant of the atmospheric mixing angle. In the presence of quantum entanglement, CPV sensitivity was reduced from $8.50\sigma$ to $8.25\sigma$ and $7.19\sigma$, respectively, for $\delta_{CP}^{\text{true}}$ values between $0^{\circ}$ and $180^{\circ}$. In the range $\delta_{CP}^{\text{true}}=[180^{\circ}–360^{\circ}]$, the highest CPV sensitivity corresponds to the “global” minimum concurrence point with a value of $7.33\sigma$, while the “local” and NuFIT data points yield maximum sensitivities of $6.37\sigma$ and $7.04\sigma$, respectively. In contrast to CPV, the sensitivity to the mass hierarchy increase when considering the concept of entanglement. However, for octant sensitivity, due to the atmospheric mixing angle approaching maximal mixing at the points of minimum entanglement, there is no significant improvement in sensitivity in the presence of entanglement.